(MS-Excel)
Ahmed has 4 boxes, Boxes contain "15 red,10 yellow,25 red,30 yellow" balls alternatively. Write excel formulas to get:

a-How many total balls he has?

b-Number of boxes he has?

c-The box having largest number of balls?

d-The box having smallest number of balls?

e-Average of balls available?
Answer #1

a) =SUM(C1:C4)

b) =COUNT(A1:A4)

c) =MAX(C1:C4)

d) =MIN(C1:C4)

e) =AVERAGE(C1:C4)